OpenStudy (anonymous):

Choose the factored form of the given equation. 6x^2 + 2x - 8 = 0

OpenStudy (anonymous):

@Peasant1

OpenStudy (anonymous):

(6x - 8)(x + 1) = 0 (6x + 1)(x - 8) = 0 (6x + 8)(x - 1) = 0

OpenStudy (anonymous):

^ My choices

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

mmk :)

OpenStudy (anonymous):

It's the last choice. :)

OpenStudy (anonymous):

mmk. And why is that?

OpenStudy (anonymous):